The answer is 47, show steps

21. How many different sums of money can you make with three pennies, a nickel, a dime, and two quarters?
Answer #1

here number of ways to select pennies =4 (as we can select 0,1,2,3 pennies)

number of ways to select a nickel=2 (as we can select 0,1 nickel)

number of ways to select a dime=2 (as we can select 0,1 dime)

number of ways to select a quarter=3 (as we can select 0,1,2 dime)

hence total number of ways to have different sums of money =4*2*2*3-1=47

( we have deducted 1 from above as there is one way where neither of pennies; nickel; dime and quarter are selected)